OpenStudy (calculusxy):

MEDAL! Which of the following is true for a circle with a circumference of about 110 meters. A. radius about 35 m B. Area about 44m^2 C. Diameter about 35 m D. Area about (17.5)^2 x 3.14m^2

OpenStudy (calculusxy):

@dan815

OpenStudy (calculusxy):

@amistre64

OpenStudy (calculusxy):

I got C

OpenStudy (calculusxy):

Is that correct?

jimthompson5910 (jim_thompson5910):

C = pi*d 110 = 3.14*d 110/3.14 = d d = 35.0318 so yeah the diameter is roughly 35 meters. There is one more factoid that's true as well.